OpenStudy (anonymous):

Help! Find an equation of the line that contains the given point and is parallel to the given line. (2,3), y=2x-7

OpenStudy (zpupster):

y = mx+b y= 2x-7 so m=2 parallel lines have identical slopes point-slope form y-y1=m(x-x1) we are given (x1,y1) as (2,3) just plug them in, and your slope is 2